Evaluate (g•7)(7) given that f(x) = Vx+2 and g(x) = x? - 5.
Answer #1

solution f(x) = √x+2 g(x) = x²-5 (gof)(7) = g(7)• 4(7) (57+2). (7-5) 9.(44) 3.uy (gof)(7) 132 (804) (7) - g CF(71) = g( 5x+7)